OpenStudy (anonymous):

i have 1 more that is long. hopefully u will remember the one with the cruise ships. same thing. caribbean paradise has 28 more staterooms that the pacific paradise. the mediteranean paradise has 32 fewer staterooms than 3 times the number of staterooms on the pacific. find the number of staterooms for each ship if total staterooms for the 3 ships is 976.

OpenStudy (anonymous):

C=P+28 M=3P-32 C+M+P=976 (P+28)+(3P-32)+P=976 5P-4=976 5P=980 P=196 sub back into first equations: C=224 M=475

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

i submitted my answers and it said that M=556 not 475.

OpenStudy (anonymous):

Sorry, must have made a calculation error there somewhere.

OpenStudy (anonymous):

no prob. it happens. at least you all help me pass the review test. thanks again. will be back here on monday night for the test.